# Learning Confidence Bounds for Classification with Imbalanced Data

### v1

@@ 1 Introduction
In a supervised classification setting, class imbalance refers to scenarios where the proportion for each class in a given dataset is uneven. Real world datasets spanning many domains such as medicine, business, and bio-informatics contain some degree of imbalance [18]. For example, in a medical dataset depicting a rare disease diagnosis, there will inherently be more examples of healthy individuals than individuals with a positive diagnosis of the disease. In this case the few examples of the positive diagnosis are the minority class and the healthy individuals are in the majority class. Even though there are fewer examples of the minority class, there is a high importance to correctly classify a positive instance of the disease since a delay in treatment will incur serious ramifications.
@@
Undersampling suffers from the issue of wasting useful information in the disregarded samples. Oversampling produces a change in the data distribution which can lead to undesirable classification behaviour. Cost-sensitive algorithms which weight minority class samples with higher importance suffer from the same effects as oversampling by replication, causing the classifier to be over confident around the minority class. More involved cost-sensitive learning algorithms are often specific to a classifier and require retraining from scratch, offering little flexibility to provide a general solution to class imbalance classification.
@@ 2 Related Work
Learning from imbalanced data sources is an area that has attracted a lot of interest over the years [4, 20, 24]. However, most approaches to deal with class imbalance rely on addressing oversimplified versions of the problem. Nevertheless, these are popular amongst practitioners as, without intervention, standard learning algorithms lead to classifiers which perform poorly on minority classes since the focus during training is skewed towards the majority class in order to obtain the lowest average loss.
@@
One of the most popular techniques when dealing with imbalance, combining undersampling the majority class with oversampling the minority class, is Synthetic Minority Oversampling Technique (SMOTE) [8]. The oversampling procedure interpolates between existing minority class data points in order to synthesise new examples. The motivation for this is that it will create a classifier with a broader decision region for the minority class and alleviate the issues when oversampling by replication. However, in some cases SMOTE has been shown to not conform to the original data distribution well, which can introduce undesirable patterns and biases to the data [16].
@@
Over and undersampling based methods occlude the imbalance information when transforming an unbalanced dataset into balanced. However, this knowledge of imbalance is informative, since having fewer samples in the minority class informs us that there is more uncertainty associated with the class’s distribution. The classifier should take this into account when assigning confidence around the classification of the minority class.
@@
An intuitive approach, inspired by importance sampling [25], is to change the weighting or ‘importance’ of data points. The simplest weighting to use is the inverse proportion a class’s occurrence. These weightings are then applied in the loss function for a Neural Network (NN) [13, 22], logistic regression [10, 26] or Support Vector Machine (SVM) [38].
More involved cost-sensitive learning algorithms exist for SVMs [6], NNs [5, 7, 33] and Decision Trees [11]. However, the issue with dealing with class imbalance during training is that the cost-sensitive learning algorithms are often classifier specific. Therefore, unlike the data adjustment methods which can be used with any classifier, the during training methods do not generalise. This makes them unsuitable as a general solution to class imbalance problems.
@@ 4 Method
In this work we focus on the role of the bias term $b$ and how to determine its value so that we can address class imbalance in a principled manner. For example, in the case of the minimum distance classifier, $b$ is set halfway between the class means. Similarly, for a vanilla SVM it is set halfway between the support vectors. Here, we introduce a method of setting $b$ in a position which is aware of class imbalance and class generalisation by utilising the concentration inequality in Section 3, which remains true for the projected space $\langle\phi(x),w\rangle$ due to the linearity of the inner product.
@@ 4.2 Relaxing the Constraints
Inspired by soft-margin SVMs [12], let us introduce a slack variable, $\xi_{n}$, per instance in the training set, where $\xi_{n}\in\{0,1\}$ defined such that $\xi_{n}=0$ for $x_{i}$ that contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$, and $\xi_{n}=1$ for $x_{i}$ which are allowed to exist outside of the supports of the data and do not contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$.
@@
The optimisation of $L_{sv}$ is less straightforward than $L$. We iteratively set $\xi_{n}=1$ for class support instances, as these are the points directly contributing to $\hat{\bar{R}}_{i}$. Instances from each class are selected such that $\sum_{n\in\{1,...,N\}}\xi_{n}$ for each class matches the class imbalance present in the training set. The final method is described in pseudocode in Algorithm 1.
@@ 5 Experiments
In this section we evaluate our method against popular class balancing techniques from the literature. Firstly, we show an illustrative example on a simple synthetic dataset so that the reader can solidify their understanding of our method. After this we showcase our method on commonplace datasets used for bench-marking class imbalance methods.
@@
Our experiments make use of the imbalance-learn package [29], costcla  [^1] https://albahnsen.github.io/CostSensitiveClassification/, scikit-learn [32]. Unless specified otherwise, all experiments use the default parameters from each package.
Since we use datasets with a small number of minority class points, the sample taken of the train/test splits can significantly skew the results. To gain unbiased results, we take the mean and standard deviation of 10 randomly sampled train and test sets for the evaluation of each dataset.
All of the experiments can be reproduced using our open source python package  [^2] https://github.com/mattclifford1/Confidence-Bounds-ECAI24. Our package provides practitioners with the ability to apply our class balancing method to their own data and classifiers with minimal code adaptation from a standard machine learning pipeline. All experiments in this paper are run on an AMD Ryzen 9 3900X CPU.
@@ 5.1 Synthetic Example
We train a Logistic Regression classifier [32]. We use standard (not cost-sensitive) learning algorithm on the original data as well as and SMOTE [8] adjusted training data. Figure 3 shows the training data and decision surface for the baseline classifier (trained on the original data) and SMOTE. We additionally train in a cost sensitive manner using balanced weights (BW).
@@ Results
Figure 5 shows the projected space after a minimum of Equation 12 is found. Finding the solution is relatively fast, taking 4 seconds on our 12 core CPU. We can see that $\hat{R}_{2}$ is bigger than $\hat{R}_{1}$, this is due to the higher levels of uncertainty associated with the few samples present in the minority class. This extra ‘space’ given to the minority class gives a level of caution need for good generalisation when the minority class is scarce.
@@ 5.2 Class Imbalanced Datasets
We evaluate our method on medical datasets, as most clinical conditions analysed in these affect small portions of the overall population, making them particularly relevant for our interests. Here, we choose those frequently used in the class imbalanced literature [8, 35]. For Pima Diabetes and Breast Cancer datasets we randomly sample a training set which contains a higher level of class imbalance (10:1) to fully showcase our method’s capabilities, this comes with the benefit of providing enough minority class samples to create a roughly balanced test set.
Table 2 shows the details of the datasets. We scale all dataset features to $[-1,1]$. For the Hepatitis dataset we remove the features ‘Protime’, ‘Alkphosphate’ and ‘Albumin’ 18 instances due to missing values. For the Heart disease dataset, we remove the features ‘ca’ and ‘thal’, as they contain missing values. To create an unbalanced dataset, we combine the heart disease types 3 and 4 as the positive class. For the MIMIC-III dataset [23], we perform ICU negative outcome prediction, using the data pre-processing outlined in [31].
@@ Results
Our Method .$\textbf{.673}\pm.03$$\textbf{.651}\pm.05$$\textbf{.611}\pm.08$
@@
Our Method .$\textbf{.944}\pm.02$$\textbf{.943}\pm.02$$\textbf{.946}\pm.02$
@@
Our Method .$778\pm.09$$\textbf{.716}\pm.05$$\textbf{.539}\pm.08$
@@
MIMIC ICU Baseline .$861\pm.00$$370\pm.01$$181\pm.01$
SMOTE [8] .$832\pm.00$$401\pm.01$$183\pm.01$
BW .$852\pm.00$$375\pm.02$$177\pm.01$
BMR [2] .$\textbf{.872}\pm.01$$322\pm.03$$152\pm.01$
Thresh [35] .$830\pm.02$$413\pm.05$$190\pm.02$
Our Method .$836\pm.02$$\textbf{.413}\pm.02$$\textbf{.196}\pm.01$
@@ 6 Discussion and Conclusion
In this paper we propose a novel and theoretically-principled method of tackling class imbalance by tailoring pre-trained classifiers using confidence bounds. Importantly, our approach makes fewer assumptions on the underlying data distribution than alternative resampling methods, whilst also not being specific to a single classifier – like many of the cost-sensitive learning algorithms mentioned in Section 2.
@@ 6.1 Limitations
We would like to thank Jeff Clark and Chris McWilliams for their help with the evaluation on the MIMIC III dataset. This work is supported by the UKRI Centre for Doctoral Training in Interactive AI EP/S022937/1 and UKRI Turing AI Fellowship EP/V024817/1.

### v2

@@ 1 Introduction
In a supervised classification setting, class imbalance refers to scenarios where the proportion for each class in a given dataset is uneven. Real world datasets spanning many domains such as medicine, business, and bio-informatics contain some degree of imbalance [20]. For example, in a medical dataset depicting a rare disease diagnosis, there will inherently be more examples of healthy individuals than individuals with a positive diagnosis of the disease. In this case, the few examples of the positive diagnosis are the minority class and the healthy individuals are in the majority class. Even though there are fewer examples of the minority class, there is a high importance to correctly classify a positive instance of the disease since a delay in treatment will incur serious ramifications.
@@
Undersampling removes data points from the majority class which wastes useful information in the disregarded samples. Oversampling creates additional minority class data points. This can produce a change in the data distribution which can lead to undesirable classification behaviour. Cost-sensitive algorithms which weight minority class samples with higher importance suffer from the same effects as oversampling by replication, causing the classifier to be over confident around the minority class. More involved cost-sensitive learning algorithms are often specific to a classifier and require retraining from scratch, offering little flexibility to provide a general solution to class imbalance classification.
@@ 2 Related Work
Learning from imbalanced data sources is an area that has attracted a lot of interest over the years [4, 22, 26]. However, most approaches to deal with class imbalance rely on addressing oversimplified versions of the problem. Nevertheless, these are popular amongst practitioners as, without intervention, standard learning algorithms lead to classifiers which perform poorly on minority classes since in order to minimise the average loss, the focus during training is skewed towards the majority class.
@@
One of the most popular techniques when dealing with imbalance, combining undersampling the majority class with oversampling the minority class, is Synthetic Minority Oversampling Technique (SMOTE) [8]. The oversampling procedure interpolates between existing minority class data points in order to synthesise new examples. The motivation for this is that it will create a classifier with a broader decision region for the minority class and alleviate the issues when oversampling by replication. However, in some cases SMOTE has been shown to not respect the original data distribution, which can introduce undesirable patterns and biases to the data [18].
@@
Over and undersampling based methods occlude the imbalance information when transforming an unbalanced dataset into balanced. However, this knowledge of imbalance is informative, since having fewer samples in the minority class informs us that there is more uncertainty associated with the distribution of that class. The classifier should take this into account when assigning confidence around the classification of the minority class.
@@
An intuitive approach, inspired by importance sampling [27], is to change the weighting or ‘importance’ of the loss at certain data points. The simplest weighting to use is the inverse proportion a class’s occurrence. These weightings are then applied in the loss function for a Neural Network (NN) [15, 24], logistic regression [11, 28] or Support Vector Machine (SVM) [40].
More involved cost-sensitive learning algorithms exist for SVMs [6], NNs [5, 7, 35] and Decision Trees [13]. However, the issue with dealing with class imbalance during training is that the cost-sensitive learning algorithms are often classifier specific. Therefore, unlike the data adjustment methods which can be used with any classifier, the methods used during training do not generalise. This makes them unsuitable as a general solution to class imbalance problems.
@@ 4 Method
In this work we focus on the role of the bias term $b$ and how to determine its value so that we can address class imbalance in a principled manner. For example, in the case of the minimum distance classifier, $b$ is set halfway between the class means. Similarly, for a vanilla SVM it is set halfway between the support vectors. Here, we introduce a method of setting $b$ in a position which is aware of class imbalance and class generalisation by utilising the concentration inequality in Eq. 3.2, which remains true for the projected space $\langle\phi(x),w\rangle$ due to the linearity of the inner product.
@@ 4.2 Relaxing the Constraints
Inspired by soft-margin SVMs [14], let us introduce a binary slack variable, $\xi_{n}$, per instance in the training set, where $\xi_{n}\in\{0,1\}$ defined such that $\xi_{n}=0$ for $x_{i}$ that contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$, and $\xi_{n}=1$ for $x_{i}$ which are allowed to exist outside of the supports of the data and do not contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$.
@@
The optimisation of $L_{sv}$ is less straightforward than $L$. We iteratively set $\xi_{n}=1$ for class support instances, as these are the points directly contributing to $\hat{\bar{R}}_{i}$. Instances from each class are selected such that $\sum_{n\in\{1,...,N\}}\xi_{n}$ for each class matches the class imbalance present in the training set. The final method is described in pseudocode in Algorithm 1. We provide an alternative approach using continuous slack variables in Appendix A.
@@ 5 Experiments
In this section we evaluate our method against popular class balancing techniques from the literature. First, we show an illustrative example on a simple synthetic dataset so that the reader can solidify their understanding of our method. After this, we showcase our method on commonplace datasets used for bench-marking class imbalance methods.
@@
Our experiments make use of the imbalance-learn package [31], costcla [12], scikit-learn [34]. Unless specified otherwise, all experiments use the default parameters from each package.
Since we use datasets with a small number of minority class points, the train/test splits can significantly skew the results. To gain unbiased results, we take the mean and standard deviation of 10 randomly sampled train and test sets for the evaluation of each dataset.
All of the experiments can be reproduced using our open source python package [10]. Our package provides practitioners with the ability to apply our class balancing method to their own data and classifiers with minimal code adaptation from a standard machine learning pipeline. All experiments in this paper are run on an AMD Ryzen 9 5950X CPU.
@@ 5.1 Synthetic Example
We train a Logistic Regression classifier [34]. We use a standard (not cost-sensitive) learning algorithm on the original data and using SMOTE [8] adjusted training data. Figure 3 shows the training data and decision surface for the baseline classifier (trained on the original data) and SMOTE. We additionally train in a cost sensitive manner using balanced weights (BW).
@@ Results
Figure 5 shows the projected space after a minimum of Equation 12 is found. Finding the solution is relatively fast, taking 3 seconds on our 16 core CPU. We can see that $\hat{R}_{2}$ is bigger than $\hat{R}_{1}$, this is due to the higher levels of uncertainty associated with the few samples present in the minority class. This extra ‘space’ given to the minority class gives a level of caution needed for good generalisation when the minority class is scarce.
@@ 5.2 Class Imbalanced Datasets
We evaluate our method on medical datasets, as most clinical conditions analysed in these affect small portions of the overall population, making them particularly interesting. Here, we choose those frequently used in the class imbalanced literature [8, 37]. For Pima Diabetes and Breast Cancer datasets we randomly sample a training set which contains a higher level of class imbalance (10:1) to fully showcase our method’s capabilities, this comes with the benefit of providing enough minority class samples to create a roughly balanced test set.
Table 2 shows the details of the datasets. We scale all dataset features to $[-1,1]$. For the Hepatitis dataset we remove the features ‘Protime’, ‘Alkphosphate’ and ‘Albumin’ and 18 instances due to missing values. For the Heart disease dataset, we remove the features ‘ca’ and ‘thal’, as they contain missing values. To create an unbalanced dataset, we combine the heart disease types 3 and 4 as the positive class. For the MIMIC-III dataset [25], we perform ICU negative outcome prediction, using the data pre-processing outlined in [33].
@@ Results
Our Method .$\textbf{.672}\pm.03$$\textbf{.651}\pm.05$$\textbf{.611}\pm.08$
@@
Our Method .$\textbf{.943}\pm.02$$\textbf{.942}\pm.02$$\textbf{.946}\pm.02$
@@
Our Method .$775\pm.09$$\textbf{.714}\pm.05$$\textbf{.535}\pm.08$
@@
MIMIC ICU Baseline .$841\pm.01$$429\pm.03$$212\pm.02$
SMOTE [8] .$822\pm.01$$431\pm.02$$199\pm.01$
BW .$839\pm.01$$410\pm.02$$195\pm.02$
BMR [2] .$\textbf{.852}\pm.01$$405\pm.04$$2\pm.02$02
Thresh [37] .$844\pm.01$$421\pm.04$$208\pm.02$
Our Method .$802\pm.02$$\textbf{.477}\pm.04$$\textbf{.221}\pm.02$
@@ 6 Discussion and Conclusion
In this paper we propose a novel and theoretically-principled method of tackling class imbalance by tailoring pre-trained classifiers using confidence bounds. Importantly, our approach makes fewer assumptions on the underlying data distribution than alternative resampling methods, whilst not being specific to a single classifier – like many of the cost-sensitive learning algorithms mentioned in Section 2.
@@ 6.1 Limitations
We would like to thank Jeff Clark and Chris McWilliams for their help with the evaluation on the MIMIC III dataset. This work is supported by the UKRI Centre for Doctoral Training in Interactive AI EP/S022937/1, UKRI Turing AI Fellowship EP/V024817/1 and the EU H2020 TAILOR, GA No 952215.
Appendix A Continuous Slack Variables
An alternative to the binary slack variables, $\xi_{n}$, presented in Sec. 4.2 is to use continuous values such that $\xi_{n}=0$ for $x_{i}$ that contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$, and $\xi_{n}=||x_{n}-\bar{\phi}_{S_{i}}||$ for $x_{i}$ which are allowed to exist outside of the supports of the data and do not contribute towards $\hat{\bar{R}}_{i}$ and $N_{i}$. This definition of slack variables is more similar to how they are defined in soft-margin SVMs [14], but adds additional complexity when compared to their binary counterpart.
Algorithm 1 remains unchanged, apart from the new definition of the new continuous slack variables. Our code provides the option for a user to choose either binary or continuous slack variables [10].
Table 4 shows the results on the datasets from Tab. 2 using continuous slack variables. The scores remain similar to those with the binary slack variables shown in Tab. 3 across all datasets.
Table 4: Evaluation of all medical datasets with continuous slacks.
Methods Accuracy G-Mean F1
Pima Diabetes Baseline .$540\pm.04$$181\pm.23$$137\pm.18$
SMOTE [8] .$595\pm.01$$495\pm.03$$396\pm.04$
BW .$585\pm.02$$476\pm.06$$373\pm.08$
BMR [2] .$661\pm.02$$628\pm.04$$575\pm.05$
Thresh [37] .$661\pm.02$$628\pm.04$$575\pm.05$
Our Method .$\textbf{.672}\pm.03$$\textbf{.651}\pm.05$$\textbf{.611}\pm.08$
Breast Cancer Baseline .$917\pm.02$$917\pm.02$$913\pm.02$
SMOTE [8] .$793\pm.09$$768\pm.12$$737\pm.14$
BW .$921\pm.02$$921\pm.02$$918\pm.02$
BMR [2] .$887\pm.08$$880\pm.09$$869\pm.11$
Thresh [37] .$935\pm.06$$932\pm.07$$929\pm.09$
Our Method .$\textbf{.943}\pm.02$$\textbf{.941}\pm.02$$\textbf{.945}\pm.02$
Hepatitis Baseline .$\textbf{.827}\pm.06$$623\pm.15$$485\pm.16$
SMOTE [8] .$804\pm.02$$299\pm.31$$224\pm.24$
BW .$768\pm.09$$681\pm.09$$499\pm.10$
BMR [2] .$804\pm.05$$656\pm.11$$493\pm.10$
Thresh [37] .$786\pm.08$$686\pm.09$$516\pm.09$
Our Method .$775\pm.09$$\textbf{.714}\pm.06$$\textbf{.535}\pm.08$
Heart Disease Baseline .$\textbf{.885}\pm.02$$784\pm.04$$718\pm.05$
SMOTE [8] .$844\pm.03$$700\pm.07$$602\pm.09$
BW .$842\pm.02$$781\pm.06$$664\pm.06$
BMR [2] .$4.8$04 .$819\pm.06$$704\pm.07$
Thresh [37] .$837\pm.05$$829\pm.03$$703\pm.06$
Our Method .$878\pm.02$$\textbf{.835}\pm.05$$\textbf{.740}\pm.06$
MIMIC ICU Baseline .$841\pm.01$$429\pm.03$$212\pm.02$
SMOTE [8] .$822\pm.01$$431\pm.02$$199\pm.01$
BW .$839\pm.01$$410\pm.02$$195\pm.02$
BMR [2] .$\textbf{.852}\pm.01$$405\pm.04$$2\pm.02$02
Thresh [37] .$844\pm.01$$421\pm.04$$208\pm.02$
Our Method .$803\pm.02$$\textbf{.476}\pm.04$$\textbf{.220}\pm.02$